OpenStudy (anonymous):

BEST RESPONSE WILL GET A MEDAL AND A FAN For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 4x − y = 8 2x + 3y = 9 8 − 4x 8 + 4x −8 + 4x −8 − 4x

OpenStudy (anonymous):

help please

OpenStudy (jiteshmeghwal9):

4x-y=8 Add y both sides wht do u gt ?

OpenStudy (anonymous):

i dont understand

OpenStudy (jiteshmeghwal9):

Adding y both sides 4x-y+y=8+y 4x=8+y

OpenStudy (jiteshmeghwal9):

Gt it till right here ?

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

what do i do next

OpenStudy (jamesj):

Actually what you are being asked to do is find an expression for y in terms of x, and substitute that into the second equation. So if your two equations are 4x − y = 8 2x + 3y = 9 Then from the first equation we have that -y = 8 - 4x hence y = -8 + 4x or y= 4x - 8 This is the expression you would substitute into the second equation, as in 2x + 3y = 9 thus 2x + 3(4x - 8) = 9 Note the the question is NOT asking you to solve the equations. It is just asking you to find an expression for y from the first of each pair of equations which you would substitute into the second equation.

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

so it's -8-4x?

OpenStudy (anonymous):

hello

OpenStudy (jamesj):

No 4x - y = 8 Hence -y = 8 - 4x and thus y = -(8-4x) y = 4x - 8

OpenStudy (anonymous):

but what if that isnt one of the choices

OpenStudy (jamesj):

But it IS one of the choices.

OpenStudy (jamesj):

4x - 8 is equal to exactly one of these expressions: 8 − 4x 8 + 4x −8 + 4x −8 − 4x

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

8-4x?

OpenStudy (jamesj):

4x - 8 Can it be 8 - 4x, I don't think so, because if it were then 4x - 8 = 8 - 4x = -(4x - 8)

OpenStudy (jamesj):

Remember that a + b = b + a. This is called the commutative property of addition. Hence 4x - 8 = -8 + 4x